OpenStudy (anonymous):

Hi, everyone. I have this mathematical problem to prove: sin(a)+sin(b)+sin[(a+b)/2] = 4cos(a/2)cos(b/2)sin[(a+b)/2]

OpenStudy (zehanz):

Looking at the LEFT side of the thing, you can add two sines by using Simpson's formula: sin(a)+sin(b)=2sin[(a+b)/2] IMO that looks promising ;) Now I'll look at the right side...

OpenStudy (anonymous):

thanks :)

OpenStudy (zehanz):

Sorry, that Simpson formaula didn't came out right: sin(a)+sin(b)=2sin[(a+b)/2]cos[(a-b)/2]

OpenStudy (anonymous):

oh, ok i'll re-write that :)

OpenStudy (zehanz):

So now I've got on the left hand side: 2sin[(a+b)/2]cos[(a-b)/2]+sin[(a+b)/2] Now sin[(a+b)/2] can be factored out: sin[(a+b)/2](2cos[(a-b)/2]+1)

OpenStudy (anonymous):

where did your 2come from 2sin[(a+b)/2] and when you move it to the right it would be minus

OpenStudy (zehanz):

The sin[(a+b)/2] is also a factor on the right hand side, so now we have to prove: (2cos[(a-b)/2]+1=4cos(a/2)cos(b/2) Do you see what I mean?

OpenStudy (anonymous):

I see it :)

OpenStudy (zehanz):

OK, let's focus on 2cos[(a-b)/2]+1 You probably know the formula cos(a-b)=cos(a)cos(b)+sin(a)sin(b), so: \[2\cos \frac{ a-b }{ 2 }+1=2\cos \left( \frac{ a }{ 2 }-\frac{ b }{ 2 } \right)+1=2\cos \frac{ a }{ 2 }\cos \frac{ b }{ 2 }+2\sin \frac{ a }{ 2}\sin \frac{ b }{ 2 }+1\]

OpenStudy (anonymous):

Ok, i get it. 'till there, what should i do next?

OpenStudy (zehanz):

Look at the right hand side. Only \[4\cos \frac{ a }{ 2 }\cos \frac{ b }{ 2 }\]left there, remember? Now it looks a lot like what we have found on the left hand side, in fact it is\[2\cos \frac{ a }{ 2 }\cos \frac{ b }{ 2 }+2\cos \frac{ a }{ 2 }\cos \frac{ b }{ 2 }\]So if I compare these two sides, I think I have to prove that\[2\sin \frac{ a }{ 2 }\sin \frac{ b }{ 2 }+1=2\cos \frac{ a }{ 2 }\cos \frac{ b }{ 2 }\]

OpenStudy (zehanz):

Looking at this, I'm thinking that this won't hold for every a and b, I'm afraid...

OpenStudy (anonymous):

ok, but it's fine ;) thx a lot!

OpenStudy (zehanz):

YW!